Grade 11Analytical Geometry

Prove that-
If a b c are 3 points on hyperbola (xy=c^2) such that ab subtends right angle at c. then ab is parallel to the normal to hyperbola at c.

Disha
7 Years agoGrade 11

1 Answer

Rajat
7 Years ago
Let the coordinates of A,B,C be (ct1,c/t1), (ct2,c/t2), (ct3,c/t3)
slope of CA= -1/(t1t3)
slope of CB= -1/(t2t3) 
(-1/(t1t3))(-1/(t2t3))=(-1) (since CA is perpe. to CB)
or, (-1/(t32))(-1/t1t2)=(-1)
or, (slope of tangent at point C )(slope of AB)=(-1)
so, tangent at C is perpendicular to AB, thus the normal at C is parallel to AB.
PROVED!